OpenStudy (anonymous):

What is the slope of the line that passes through (-3, 3) and (-3, 6)? A) -1/2 B) 0 C) 1/2 D) Undefined

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

so, \[\frac{ 6 }{ -3 } - \frac{ 3 }{ -3 }\] which equals \[\frac{ 3 }{ 0 }\]

OpenStudy (anonymous):

i think this is undefined

OpenStudy (anonymous):

why do you say that? @needhelpexplainplz

OpenStudy (anonymous):

because it makes a vertical line

OpenStudy (anonymous):

Oh, is it my work at the top correct though?

OpenStudy (anonymous):

yes i believe so

OpenStudy (abb0t):

A vertical line has "undefined slope" because all points on the line have the same x-coordinate.

OpenStudy (anonymous):

you can close your question now

OpenStudy (anonymous):

I know..but i can't do that whenever i still don't get the question. I don't want the answer just like you and @abb0t gave me. I want an explanation to why it is that.

OpenStudy (anonymous):

@Deadly_Roses @RCCB

OpenStudy (abb0t):

They already gave you an explantion: Slope = m = \(\sf \large \frac{\Delta y }{\Delta x} = \frac{y_2-y_1}{x_2-x_1}\) you subtract the second coordinate from the first. it is the change in coordinates.

OpenStudy (anonymous):

@abb0t Don't even mind replying back to what i just said above, i could careless about what your peasant mind has to say to me so therefore leave, thank you have a nice day(:

OpenStudy (anonymous):

@https.wild_ um excuse me i know your talking to him but you dont have a reason to be rude we shouldnt have given you the answer >:(

OpenStudy (abb0t):

I'm disappointed that you fail to understand and grasp how this problems works, despite the explanations given to you by multiple users. I suppose you need to study your elementary mathematics.

OpenStudy (anonymous):

I didn't wanted the answer..i wanted an explanation @needhelpexplainplz

OpenStudy (anonymous):

i did give you the explaination i said that it was that answer because it was a veritical line!

OpenStudy (abb0t):

No one gave you the answer here. Close the question, you're just trolling now.

OpenStudy (deadly_roses):

how to find the slope.....(this is on a different problem, but just to show you how) To find the slope, we will need two points from the line. Pick two x's and solve for each corresponding y: If, say, x = 3, then y = ( 2/3 )(3) – 4 = 2 – 4 = –2. If, say, x = 9, then y = ( 2/3 )(9) – 4 = 6 – 4 = 2. (By the way, I picked the x-values to be multiples of three because of the fraction.

OpenStudy (anonymous):

Thank you, that makes more sense to me. @Deadly_Roses

OpenStudy (deadly_roses):

your welcome :)